OpenStudy (anonymous):

Assume the triangle has the given measurements. Solve for the remaining sides and angles a) C=100°, a=(1/2),b=(1/3) b) A=(π/6,a=17.4,b=19.6

OpenStudy (anonymous):

@Jack1

OpenStudy (jack1):



OpenStudy (jack1):

so do u recon for (a) we use sine rule or cosine rule...?

OpenStudy (anonymous):

cosine

OpenStudy (jack1):

coolski, me to so \(\large c^2 = a^2 + b^2 - 2ab(\cos (C))\) solve for c = ...? @washcaps ?

OpenStudy (anonymous):

ok hold on

OpenStudy (jack1):

then use Sine rule to solve for other angles... easier that way ;)

OpenStudy (anonymous):

c=.65

OpenStudy (jack1):

i'll just work it out too, sorry man, 1 sec

OpenStudy (jack1):

0.647... so 0.65, so perfect man

OpenStudy (jack1):

now use sine rule to work out other angles

OpenStudy (anonymous):

so sinB/1/3? sinA/1/2?

OpenStudy (jack1):

\[\Large \frac a{\sin A} = \frac c{\sin C}\]... solve for A first.. then do the same for B

OpenStudy (jack1):

ah, ur way is easier solving for angle, sorry man

OpenStudy (jack1):

\[\Large \frac {\sin A}a = \frac {\sin C}c\]

OpenStudy (anonymous):

my calculator is screwing up. I know the equation is sinA/1/2 = sin(100)/0.65. Could you solve?

OpenStudy (jack1):

sure sin (100)/0.647297 = 1.52142 = Sin A/0.5 so Sin A = 0.76071 so A = 49.52 degrees

OpenStudy (jack1):

180 - 100 - 49.52 = B = 30.48 ish...?

OpenStudy (anonymous):

real quick this website http://www.calculatorsoup.com/calculators/geometry-plane/triangle-law-of-sines.php says that A=49.24?

OpenStudy (jack1):

maybe i rounded too early...?

OpenStudy (anonymous):

oh ok so what would be the right answer? the website or yours?

OpenStudy (jack1):

maybe u used 0.65 instead of 0.647297 on the website...?

OpenStudy (anonymous):

oh I did sorry. Yours is correct then

OpenStudy (jack1):

cools ;) either way they're both pretty close

OpenStudy (jack1):

k, u got this man, imma finish my assignment, slaters ;)

OpenStudy (anonymous):

Thanks. I could use your help later tonight if your on haha :)

OpenStudy (jack1):

prob not, its 3:30 am here, will be sleepin like the dead tomorro night ;)

OpenStudy (anonymous):

not on this. different question lol oh ok man thanks though

OpenStudy (jack1):

sorry man :(

OpenStudy (jack1):

np night